There are 400 kg/min of water being handled by a pump. The lift is from a 20-m deep well and the delivery velocity is 15 m/s. Find (a) the change of potential energy in kJ/kg, (b) the kinetic energy in kJ/kg, and (c) the required power of the pumping unit in kW. Assume g = 9.75 m/s2.
1
Expert's answer
2020-10-08T15:11:03-0400

In one minute: 400 kg of water is lifted 20 m and accelerated to a speed of 15 m/s (presumably from rest).

a. Change in PE = mgh = "400\u00d79.75\u00d720 = 78000 J"

b. Change in KE = "\frac{1}{2}mv^2 = \frac{400}{2} × 15^2 = 45000 J"

c. Total energy imparted to water per minute = 123 kJ

Thus required power = 123 / 60 = 2.05 kW

This assumes that we neglect all resistances.